OpenStudy (anonymous):

how do you find the recursive formula for 0 ,10, 15, 17.5, 18.75, ...

OpenStudy (anonymous):

first amount = 0 second amount increases in 10 third amount increases in half of 10, then 10 + 5 fourth amount increases in half of 5, then 15 + 2.5 fifth amount increases in half of 2.5, then 17.5 + 1.25 you have: 0 10 10 + 10/2 15 + 10/4 17.5 +10/8 Can you figure out? \[a _{0} = 0\] \[a _{n} = a _{n-1} + \frac{ 10 }{ \left( 2 \right)^{n-1}}\]